a promoter needs to make $93,500 from a sale of 3900 tickets . SHe charges $22.00 for some tickets and $33.00 for others. If there is x amoount of the $22.00 tickets and y of the $33.oo tickets sold, write an equation that states the sum of the tickets sold is $3900
Answer by richwmiller(17219)

x+y=3900,
22*x+33*y=93500
x=3900-y
22*(3900-y)+33*y=93500
85800-22y+33*y=93500
11*y=7700
y=700
x=3900-y
x=3200 at 22 and y=700 at 33
check
3200+700=3900
22*x+33*y=93500
22*3200+33*700=93500
70400+23100=93500
93500=93500
ok